Solve the system by substitution.
x + 3y = –8
5x + 2y = 12

Answer by checkley75(3666)

X+3Y=-8, X=-3Y-8 NOW SUBSTITUTE (-3Y-8) FOR X & SOLVE FOR Y IN THE SECOND EQUATION.
5(-3Y-8)+2Y=12
-15Y-40+2Y=12
-13Y=12+40
-13Y=52
Y=52/-13
Y=-4 ANSWER.
X+3*-4=-8
X-12=-8
X=-8+12
X=4 ANSWER.